OpenStudy (anonymous):

5x + 4y = -5 -5x – 7y = -10 can somebody help me with that question ?

OpenStudy (mertsj):

Add the two equations. Did you do that?

OpenStudy (anonymous):

\[\left( 5x +4y \right)+\left( -5x -7y \right)=-5+\left( -10 \right)\] \[4y -7y =-15\] \[-3y =-15\] \[y =5\] \[x =-5\]